User name templon
Log entry time 17:55:08 on September28,2000
Entry number 48446
1) aerogel cerenkov cabled up in right spectrometer (elektron det pkg but hadron
..... spect for e97111. aerogel right cabled to adc in slot 23, patch cable 1 goes
..... channel 48 of ADC, patch cable 16 goes to channel 63.
..... aerogel left cabled to adc in slot 24, signal (patch cable) 1 to chan 48,
..... signal 16 cabled to chan 64.
2) deadtime pulser in same arm found dangling. reconnected to tdc in slot
..... 19, channel 84
3) L arm, plane 4 VDC, cable 12 found shifted to left at TDC input. replugged.
4) L arm still has noisy VDC. Bogdan and Mark J checking.
5) Brad Sawatzky has already made initial changes to detmap.config to
..... make espace grok physical package swap plus our detpkg function swap